OpenStudy (anonymous):

Intermediate Value Theroem help: Show that there exista ln^x = e^(-x) on the interval (1,2)

OpenStudy (amistre64):

the notation is confusing

OpenStudy (amistre64):

if you plug in 1 and 2 into the "function" and it produces different signs; then it had to cross at some point in the interval

OpenStudy (amistre64):

as long as the function is continuous

OpenStudy (anonymous):

we were supposed to solve it w/out a calulator. somehow we have to know that 1/ln2 is less that e^2 so ln 2 is greater tahn 1

OpenStudy (amistre64):

ln^x = e^(-x) ; this notation makes no sense, specifically that ln^x ... whats it mean?

OpenStudy (anonymous):

really...shud it be lnx?

OpenStudy (amistre64):

.... i dunno, i aint got your material to compare with :)

OpenStudy (anonymous):

let me shwo u what we have so far...im confused about teh rest

OpenStudy (anonymous):

i guesss its lnx

OpenStudy (anonymous):

after evaluating ln(1) = ln1 - e^-1 = 0 - 1/e = - (1/e) ln(2) = ln2 - e^-2 = ln 2 - 1/e^2 1/ln 2 is less than e^2 therefore ln 2 is greater tahn 1

OpenStudy (amistre64):

"Intermediate Value Theroem" simply says that if the ends of the interval [a,b] are opposite in sign; that the function had to cross at least once between them.

OpenStudy (amistre64):

it has nothing to do with greater or less than; it has to do with sign change

OpenStudy (anonymous):

my teacher gave some hint as to\[\log_{e}2 = 1 \] so ln2 is less tahn 1

OpenStudy (amistre64):

ok, lets try out the less than assumption